HIGH RESOLUTION ANALYSIS OF THE ν_{6}, ν_{7}, ν_{8} AND ν_{9} BANDS OF H^{15}NO_{3}

F. KELLER, \underline{A. PERRIN}, J.-M. FLAUD, Laboratoire de Photophysique Moléculaire, CNRS, Université Paris Sud, Campus d’Orsay, Bat 210, 91405 Orsay Cedex, France; J.W. JOHNS, and ZHENG-FANG LU, Steacie Institute for Molecular Sciences, National Research Council of Canada, 100 Sussex Drive, Ottawa, KIA OR6, Canada; E.C. LOOI, Department of Physics, Faculty of Science, National University of Singapore, Lower Kent Ridge Road, Singapore, 119360 Singapore.

The analysis of the ν_{6}, ν_{7}, ν_{8} and ν_{9} bands of H^{15}NO_{3} located at 646.9641, 578.4719, 743.6166 and 458.2917cm^{−1} respectively has been carried out in the 400-800cm^{−1} region using high resolution Fourier transform spectra recorded at Ottawa. Using the ground state energy levels calculated from the v=0 rotational constants of H^{15}NO_{3} ^{a}, it was possible to assign the A-type ν_{6} and ν_{7} bands and the C-type ν_{8} and ν_{9} bands of H^{15}NO_{3} up to high J and K_{a} rotational quantum numbers. The v_{6}=1, v_{7}=1, v_{8}=1 and v_{9}=1 experimental energy levels were then introduced in a least squares fit calculation and precise upper state Hamiltonian constants (band centers and rotational constants) were determined allowing one to reproduce the infrared data to within the experimental uncertainty.

aA.P.Cox, M.C.Ellis, C.J.Attfield and A.C.Ferris, J. of Mol. Struct. \underline{320}, 91 (1994)